# RWZC: A Model-Driven Approach for Learning-based Robust Wyner-Ziv Coding

### v1

@@ Abstract
In this paper, a novel learning-based Wyner-Ziv coding framework is considered under a distributed image transmission scenario, where the correlated source is only available at the receiver. Unlike other learnable frameworks, our approach demonstrates robustness to non-stationary source correlation, where the overlapping information between image pairs varies. Specifically, we first model the affine relationship between correlated images and leverage this model for learnable mask generation and rate-adaptive joint source-channel coding. Moreover, we also provide a warping-prediction network to remove the distortion from channel interference and affine transform. Intuitively, the observed performance improvement is largely due to focusing on the simple geometric relationship, rather than the complex joint distribution between the sources. Numerical results show that our framework achieves a $1.5$ dB gain in PSNR and a $0.2$ improvement in MS-SSIM, along with a significant superiority in perceptual metrics, compared to state-of-the-art methods when applied to real-world samples with non-stationary correlations.
@@ I Introduction
Research on DSC dates back to the 1970s, when Slepian and Wolf [1] established the optimal rate region for independently encoding two correlated sources (SW problem). This was followed by the study of the Wyner-Ziv (WZ) problem (source coding with decoder-only side information) [2], the multi-terminal (MT) problem (source coding for correlated multiple users) [3], and the chief executive officer (CEO) problem (source coding for correlated sources with a common remote source) [4]. These foundamental works provide key theoretical insights that have guided the design of compression and transmission systems in real-world applications. When combining DSC with transmission, the integration of source and channel coding becomes especially critical. Joint source-channel coding (JSCC) has been shown to outperform separate source-channel coding in scenarios with limited block lengths, despite its higher implementation complexity. However, this problem has been solved by the rapid advances in deep learning (DL) technologies. In recent years, numerous convolutional neural network (CNN)-based frameworks for distributed compression and transmission have emerged, particularly for image sources [5, 6, 7, 8, 9, 10, 11, 12, 13]. For scenarios where both correlated sources require encoding (SW problem), Wang et al. [5] proposed a deep neural network solution incorporating a cross-attention module being aware of channel state information. In [6], the authors explored distributed audio-visual parsing using a multi-modal transformer and deep JSCC. In the case where side information is only available at the decoder (WZ problem), Ayzik and Avidan [8] introduced a side-information-finder block to perform plane-wise projection for deep image compression, termed the deep side information network (DSIN). Mital et al. [9, 10] further improved Ayzik’s work by integrating an entropy model and cross-attention mechanism to create a more efficient compression scheme, known as neural distributed image compression (NDIC). More recently, Yilmaz et al. [11] proposed a novel framework, DeepJSCC-WZ, which integrates an attention feature module and achieves low-latency joint source-channel coding for transmission scenarios.
@@
However, the existing frameworks for both SW and WZ problems primarily focus on highly correlated and stably distributed sources to achieve optimal compression or transmission performance, which does not accurately reflect the variability of real-world data. In scenarios where source correlation fluctuates significantly, such as image pairs with varying parallax (as shown in Fig. 1), these methods often fail to achieve satisfactory results. This performance gap can be attributed to three main reasons. First, when training on images with non-stationary correlations, instability arises during feature extraction, leading to mismatches and inaccurate representations. Second, CNN-based architectures often have insufficient receptive fields to capture the correlations in highly uncorrelated images, resulting in confusion among convolutional features and a decline in performance. Third, the non-stationary distribution of correlation in the input data significantly hinders the network’s ability to generalize across different data samples. In summary, the performance of feature-level interaction heavily relies on the stability of input distribution in data-driven architectures, and it typically performs unsatisfactorily in situations where the correlation distribution fluctuates. In such cases, a well-designed correlation model can effectively address these issues.
@@
Thus, we assume that there exists a conception, named as source state information (abbreviated by SSI), which used to describe the correlation between sources with low complexity. This approach facilitates source decorrelation, thereby preventing mismatched representations and feature confusion. To validate this conception, we consider a practical WZ scenario in this paper involving camera arrays, where two stereo cameras capture images of the same object. One camera transmits photos $\bm{x}$ to the receiver, while the receiver has access to the noise-free image $\bm{y}$ from the other camera. Note that, we assume the correlation between two images is non-stationary owing to the changeable camera poses, camera parameters or even a movable objects. To address the performance degradation caused by varying correlation, we propose three key components driven by the modeling of SSI: correlation estimation and decoupling, low-rate transmission, and SSI-aided source reconstruction. Specifically, we first estimate the SSI between the image pairs and use it for source decoupling of $\bm{x}$, where ”decoupling” refers to removing the redundant information carried by the side image $\bm{y}$. The SSI is also available at the decoder without error. Next, we apply an efficient JSCC scheme to transmit the decoupled image with limited resources. Finally, $\bm{y}$, along with SSI, aids in the reconstruction of $\bm{x}$ at the decoder.
@@
  • We propose a CNN-based autoencoder that implements low-rate JSCC, encoding the decoupled image into feature representations whose size adapts to the masked regions. This rate-adaptive mechanism utilizes non-linear transform and entropy modeling of each feature component, effectively saving the transmission resources.
  • We introduce a warping prediction reconstruction network designed to correct distortions caused by channel interference and perspective transform. An SSI-integrated mechanism is incorporated for mask prediction and distortion compensation, helping to repair pixel-level corruption in image reconstruction due to noise interference and resolution discrepancies.
@@ II-A Notation Convenience
Throughout this paper, the realization of the corresponding random variable is denoted by a lowercase letter, while a scalar is represented in normal font and a vector in bold font. For instance, we utilize the scalar $x$ and vector $\bm{x}=(x_{1},x_{2},\cdots,x_{L})$. Moreover, we use $\mathbb{R}$ and $\mathbb{C}$ to denote the sets of real numbers and complex numbers, respectively. Besides, $\bm{z}^{H}$ denotes the Hermitian transpose of matrix $\bm{z}$.
@@ II-B System Model
We consider an image transmission scenario using two stereo cameras, where the cameras are positioned at different angles to capture images of the same object, resulting in images with overlapping information. It is assumed that the camera positions are fixed. In this setup, one set of images is transmitted over a wireless channel, while another set of images, transmitted over a high-rate/error-free channel, is used at the receiver to assist in reconstructing the first set. This setup, where only side information is available at the receiver, is referred to as the WZ coding problem. Specifically, the main camera captures image $\bm{x}\in\mathbb{R}^{H\times W\times C}$ and encodes it into a $k$-length codeword $\bm{z}\in\mathbb{R}^{k}$, subject to an average power constraint: $\frac{1}{k}||\bm{z}||^{2}\leq P_{\mathrm{ow}}$. The codeword is then transmitted through an additive white Gaussian noise (AWGN) channel, which is modeled as
@@
where $\bm{n}\sim\mathcal{CN}(\bm{0},\sigma_{n}^{2}\bm{I}_{k})$ denotes the unbiased complex Gaussian noise. For further use, the metric named channel bandwidth ratio (CBR) is defined as $\frac{k}{H\times W\times C}$ to measure the resources per pixel consumed for transmission, and channel signal-to-noise ratio (SNR) is defined as $10\log_{10}(\frac{P_{\mathrm{ow}}}{\sigma_{n}^{2}})$ with unit dB. Finally, the decoder utilizes the noisy codewords to recover $\bm{x}$ with the help of side image $\bm{y}\in\mathbb{R}^{H\times W\times C}$ from the secondary camera, and finally outputs the reconstruction $\hat{\bm{x}}$.
@@
In fact, as a dual case, considering the implicit test channel $P_{\bm{Y|X}}$ existing between correlated sources, the analogy between SSI and CSI goes beyond this. For instance, both of them can be estimated apart from the encoder and fed back to optimize the transmission strategy. Besides, SSI and CSI describe source correlation and channel state distributions, respectively, but independent with the information to be transmitted itself. Similarly, we can also analogize the sources with stationary correlation distribution to block fading channels, and the sources with non-stationary correlation distribution to fast fading channels. Overall, like CSI, SSI is a characterization of system parameters, and its use in the encoder does not affect the settings of the original problem.
@@ II-C3 SSI-aided source reconstruction
where we use the decoding function $F^{-1}_{\xi}(\cdot)$ to denote the inverse operation of encoding.
@@ III Detailed Methodologies
This section presents the detailed schemes to realize the aforementioned modules. Specifically, for correlation estimation, we utilize the geometric method named as homography, which can be evaluated by existing feature matching algorithms. For decoupling method, an SSI-robust masking strategy is proposed for varying $\bm{h}$ between different image pairs, which is realized by a learnable perspective transform. Moreover, for transmission, a convolutional neural network (CNN)-based autoencoder with hyper-prior is adopted for rate adaptive transmission. Finally, for high-quality reconstruction, a distortion prediction module is proposed at the receiver for accurate warping of side image and seamless blending.
@@ III-A Homography estimation
For the homography estimation [16, 17], given an image pair, we assume there exists a matrix yielding the affine transform from $i$-th point $(p^{i}_{1},q^{i}_{1})$ in one image to point $(p^{i}_{2},q^{i}_{2})$ in the other, where $i=1,2,\cdots,S$ and $S$ denotes the number of matching points, namely
@@
which is also known as the perspective transform. Consequently the corresponding point pair $(p_{2},q_{2})$ is formulated as
@@
where $\tilde{\bm{h}}=[h_{11},h_{12},h_{13},h_{21},\cdots,h_{33}]$. This means that the homography estimation only needs four non-collinear coordinates pairs due to the 8 free parameters in $\bm{h}$. After that, this overdetermined system will be solved with the best four geometric-related pairs, which is found by algorithms. In a word, this pixel-wised affine connection uniquely determine the overlapping between image pair.
@@ III-B Homography-adaptive masking strategy
Considering the image decoupling module, we utilize a masking strategy for concealing the redundant information which has been provided by side image $\bm{y}$ at the decoder. To do this, we generate a transform-based mask for each $\bm{x}$ to remove the unnecessary information. The generation of mask $\bm{m}\in\{0,255\}^{H\times W\times C}$ can be depicted as follows: we first apply perspective transform (PT) on one image with all 255 pixel value according to $\bm{h}$, and then take logical negation operation on the transformed image, which is illustrated in Fig. 4. After that, $\bm{m}$ will conduct pixel-wise production with $\bm{x}$ to complete the masking.
@@
Figure 7: Details for JSCC codec, where $\mathrm{NLT}$ and $\mathrm{INLT}$ represent the nonlinear and inverse nonlinear transform coding, respectively, while $\mathrm{AnlT}$ and $\mathrm{SynT}$ denote the analysis transform and synthesis transform, respectively.
@@
where $\bm{x}^{*}$ and $\bm{x}_{w}$ are both functions of $\bm{x}$. Finally outputs the residual image $\bar{\bm{x}}=\bm{x}-\bm{x}^{*}$. At the same time, the learnable $\bm{h}^{*}$ will replace $\bm{h}$ and be transmitted to the decoder for high-quality reconstruction. We remark that this module functions to finetune the $\bm{h}^{*}$ to generate a learning-based mask, for solving the above problems of corrupted estimation and pixel mismatching with non-stationary data samples. Besides, the PTL-based masking strategy enjoys the advantages of negligible training overload due to the differentiable linear layer design and multi-viewpoints awareness, namely we can input multiple $\bm{h}$ into the networks for training at one time to realize a better generalization.
@@ III-C Entropy model-based transmission
For the reliable and efficient transmission, we provide a rate-adaptive JSC codec based on hyperprior-assisted nonlinear transform. The reason why rate adaption is necessary lies on the masks of irregular sizes, which yield an inefficient transmission if we send images of different masked regions with a same code rate. Fortunately, a promising solution is provided by [28] and [29] who uses the entropy model for image transmission with the assistance of latent hyperprior. Therefore, we provide a CNN-based autoencoder for rate-adaptive transmission by incorporating channel coding into the entropy model based framework. The architecture can be found in Fig. 7. Specifically, the masked image $\bar{\bm{x}}$ is first converted to features $\bm{v}$ via non-linear transform (NLT) composed of convolutional layer and normalization. After that, spatial dependencies is captured by analysis transform (ANLT) on features, and we model the feature latent as scale Gaussian variables via the hyperprior $\bm{s}$, which enables a more precise computation of entropy on each component. Then, the components of features select FC layers of different output dimensions according to the respective computed entropy, and hence realizes an effective rate adaption transmission. At the receiver, the reverse operations are conducted by synthesis transform and inverse non-linear transform modules, and finally they output a reconstructed masked image $\tilde{\bm{x}}$. For the training, the rate-distortion trade-off is utilized as the loss function, which is based on a variational inference, i.e.
@@
The detailed computation can be found in [29]. Note that $\bm{v}$ and $\bm{s}$ are both functions of $\bar{\bm{x}}$, and $a$ denotes the trade-off parameter between rate and distortions. In one word, the combination of adaptive mask generation and entropy model-based codec realizes the SSI robustness and low rate transmission. Due to above design, we find that even though RWZC encounters an unfamiliar samples during the inference process, the learning of the distribution with the homography adaption will allow the framework to effectively masking and transmitting $\bm{x}$.

### v2

@@ Abstract
In this paper, a novel learning-based Wyner-Ziv coding framework is considered under a distributed image transmission scenario, where the correlated source is only available at the receiver. Unlike other learnable frameworks, our approach demonstrates robustness to non-stationary source correlation, where the overlapping information between image pairs varies. Specifically, we first model the affine relationship between correlated images and leverage this model for learnable mask generation and rate-adaptive joint source-channel coding. Moreover, we also provide a warping-prediction network to remove the distortion from channel interference and affine transform. Intuitively, the observed performance improvement is largely due to focusing on the simple geometric relationship, rather than the complex joint distribution between the sources. Numerical results show that our framework achieves a $1.5$ dB gain in PSNR and a $0.2$ improvement in MS-SSIM, along with a significant superiority in perceptual metric, compared to state-of-the-art methods when applied to real-world samples with non-stationary correlations.
@@ I Introduction
Research on DSC dates back to the 1970s, when Slepian and Wolf [1] established the optimal rate region for independently encoding two correlated sources (SW problem). This was followed by the study of the Wyner-Ziv (WZ) problem (source coding with decoder-only side information) [2], the multi-terminal (MT) problem (source coding for correlated multiple users) [3], and the chief executive officer (CEO) problem (source coding for correlated sources with a common remote source) [4]. These foundamental works provide key theoretical insights that have guided the design of compression and transmission systems in real-world applications. When combining DSC with transmission, the integration of source and channel coding becomes especially critical. Joint source-channel coding (JSCC) has been shown to outperform separate source-channel coding in scenarios with limited block lengths, despite its higher implementation complexity. However, this problem has been solved by the rapid advances in deep learning (DL) technologies. In recent years, numerous convolutional neural network (CNN) or Transformer-based frameworks for distributed compression and transmission have emerged, particularly for image sources [5, 6, 7, 8, 9, 10, 11, 12, 13]. For scenarios where both correlated sources require encoding (SW problem), Wang et al. [5] proposed a deep neural network solution incorporating a cross-attention module being aware of channel state information. In [6], the authors explored distributed audio-visual parsing using a multi-modal Transformer and DL-based JSCC. In the case where side information is only available at the decoder (WZ problem), Ayzik and Avidan [8] introduced a side-information-finder block to perform plane-wise projection for DL-based image compression, termed the Deep Side Information Network (DSIN). Mital et al. [9, 10] further improved Ayzik’s work by integrating an entropy model and cross-attention mechanism to create a more efficient compression scheme, known as Neural Distributed Image Compression (NDIC). More recently, Yilmaz et al. [11] proposed a novel framework, DeepJSCC-WZ, which integrates an attention feature module and achieves low-latency joint source-channel coding for transmission scenarios.
@@
However, the existing frameworks for both SW and WZ problems primarily focus on highly correlated and stably distributed sources to achieve optimal compression or transmission performance, which does not accurately reflect the variability of real-world data. In scenarios where source correlation fluctuates significantly, such as image pairs with varying parallax (as shown in Fig. 1), these methods often fail to achieve satisfactory results. This performance gap can be attributed to three main reasons. First, when training on images with non-stationary correlations, instability arises during feature extraction, leading to mismatches and inaccurate representations. Second, CNN-based architectures often have insufficient receptive fields to capture the correlations in highly uncorrelated images, resulting in confusion among convolutional features and a decline in performance. Third, the non-stationary distribution of correlation in the input data significantly reduce the network’s ability to generalize across different data samples. In summary, the performance of feature-level interaction heavily relies on the stability of input distribution in data-driven architectures, thus it performs unsatisfactorily in situations where the correlation distribution fluctuates, inevitably. In such cases, a well-designed correlation model can effectively address these issues.
@@
Therefore, we assume that there exists a conception, named as source state information (abbreviated by SSI), which used to describe the correlation between sources with low complexity. This approach facilitates source decorrelation, thereby preventing mismatched representations and feature confusion. To validate the effectiveness of this conception, we consider a practical WZ scenario in this paper involving camera arrays, where two stereo cameras capture images of the same object. One camera transmits photos $\bm{x}$ to the receiver, while the receiver has access to the noise-free image $\bm{y}$ from the other camera. Note that, we assume the correlation between two images is non-stationary owing to the changeable camera poses, camera parameters or even a movable objects. To address the performance degradation caused by varying correlation, we propose three key components driven by the modeling of SSI: correlation estimation and decoupling, low-rate transmission, and SSI-aided source reconstruction. Specifically, we first estimate the SSI between the image pairs and use it for source decoupling of $\bm{x}$, where ”decoupling” refers to removing the redundant information carried by the side image $\bm{y}$. The SSI is also available at the decoder without error. Next, we apply an efficient JSCC scheme to transmit the decoupled image with limited resources. Finally, $\bm{y}$, along with SSI, aids in the reconstruction of $\bm{x}$ at the decoder.
@@
  • We propose a CNN-based autoencoder that implements low-rate JSCC, encoding the decoupled image into feature representations whose size adapts to the masked regions. This rate-adaptive mechanism utilizes non-linear transform (NLT) and entropy modeling of each feature component, effectively saving the transmission resources.
  • We introduce a warping prediction reconstruction network designed to reduce distortions caused by channel interference and perspective transform. An SSI-integrated mechanism is incorporated for mask prediction and distortion compensation, helping to repair pixel-level corruption in image reconstruction due to noise interference and resolution discrepancies.
@@ II-A Notation Convenience
Throughout this paper, the realization of the corresponding random variable is denoted by a lowercase letter, while a scalar is represented in normal font and a vector in bold font. For instance, we utilize the scalar $x$ and vector $\bm{x}=(x_{1},x_{2},\cdots,x_{L})$. Moreover, we denote the sets of real numbers and complex numbers $\mathbb{R}$ and $\mathbb{C}$, respectively. Besides, $\bm{z}^{H}$ denotes the Hermitian transpose of matrix $\bm{z}$.
@@ II-B System Model
We consider an image transmission scenario using two stereo cameras, where the cameras are positioned at different angles to capture images of the same object, resulting in images with overlapping information. It is assumed that the camera positions are fixed. In this setup, one set of images is transmitted over a wireless channel, while another set of images, transmitted over a high-rate/error-free channel, is used at the receiver to assist the reconstruction of the first set. This setup, where only side information is available at the receiver, is known as the WZ coding problem. Specifically, the main camera captures image $\bm{x}\in\mathbb{R}^{H\times W\times C}$ and encodes it into a $k$-length codeword $\bm{z}\in\mathbb{R}^{k}$, subject to an average power constraint: $\frac{1}{k}||\bm{z}||^{2}\leq P_{\mathrm{ow}}$. The codeword is then transmitted through an additive white Gaussian noise (AWGN) channel, which is modeled as
@@
where $\bm{n}\sim\mathcal{CN}(\bm{0},\sigma_{n}^{2}\bm{I}_{k})$ denotes the unbiased complex Gaussian noise. For further use, the metric named channel bandwidth ratio (CBR) is defined as $\frac{k}{H\times W\times C}$ to measure the resources per pixel consumed for the transmission, and channel signal-to-noise ratio (SNR) is defined as $10\log_{10}(\frac{P_{\mathrm{ow}}}{\sigma_{n}^{2}})$ with unit dB. Finally, the decoder utilizes the noisy codewords to recover $\bm{x}$ with the help of side image $\bm{y}\in\mathbb{R}^{H\times W\times C}$ from the secondary camera, and finally outputs the reconstruction $\hat{\bm{x}}$.
@@
In fact, as a dual case, considering the implicit test channel $P_{\bm{Y|X}}$ existing between correlated sources, the connection between SSI and CSI goes beyond this. For instance, both of them can be estimated apart from the encoder and fed back to optimize the transmission strategy. Besides, SSI and CSI describe source correlation and channel state distributions, respectively, but independent of the information to be compressed/transmitted. Similarly, we can draw an analogy between sources with stationary correlation distributions and block fading channels, while another analogy between sources with non-stationary correlation distributions and fast fading channels. Overall, like CSI, SSI is a characterization of system parameters, and its use in the encoder does not affect the settings of the original problem.
@@ II-C3 SSI-aided source reconstruction
where we use the decoding function $F^{-1}_{\xi}(\cdot)$ to represent the inverse operation of encoding.
@@ III Detailed Methodologies
This section presents the detailed schemes to realize the aforementioned modules. Specifically, for correlation estimation, we utilize the geometric method named as homography, which can be evaluated by existing feature matching algorithms. For decoupling method, an SSI-robust masking strategy is proposed for varying $\bm{h}$ between different image pairs, which is realized by a learnable perspective transform. Moreover, for transmission, a CNN-based autoencoder with hyper-prior is adopted for rate adaptive transmission. Finally, for high-quality reconstruction, a distortion prediction module is proposed at the receiver for accurate warping of side image and seamless blending.
@@ III-A Homography estimation
For the homography estimation [16, 17], given an image pair, we assume there exists a matrix yielding the affine transform between two images from one point $(p^{i}_{1},q^{i}_{1})$ to its corresponding point $(p^{i}_{2},q^{i}_{2})$, where $i=1,2,\cdots,S$ and $S$ denotes the number of all matching points, namely
@@
Consequently the corresponding point pair $(p_{2},q_{2})$ is formulated as
@@
where $\tilde{\bm{h}}=[h_{11},h_{12},h_{13},h_{21},\cdots,h_{33}]$. This means that the homography estimation only needs four non-collinear coordinates pairs due to the 8 free parameters in $\bm{h}$. After that, this overdetermined system will be solved with the best four geometric-related pairs, which is found by algorithms. In a word, this pixel-wised affine connection uniquely determine the overlapping region between image pair.
@@ III-B Homography-adaptive masking strategy
Considering the image decoupling module, we utilize a masking strategy for removing the redundant information which has been provided by side image $\bm{y}$ at the decoder. To do this, we generate a transform-based mask for each $\bm{x}$ to remove the unnecessary information. The generation of mask $\bm{m}\in\{0,255\}^{H\times W\times C}$ can be depicted as follows: we first apply PT on one image with all 255 pixel value according to $\bm{h}$, and then take logical negation operation on the transformed image, which is illustrated in Fig. 4. After that, $\bm{m}$ will conduct pixel-wise production with $\bm{x}$ to complete the masking.
@@
Figure 7: Details for JSCC codec, where $\mathrm{NLT}$ and $\mathrm{INLT}$ represent the non-linear and inverse non-linear transform coding, respectively, while $\mathrm{AnlT}$ and $\mathrm{SynT}$ denote the analysis transform and synthesis transform, respectively.
@@
where $\bm{x}^{*}$ and $\bm{x}_{w}$ are both functions of $\bm{x}$. Finally outputs the residual image $\bar{\bm{x}}=\bm{x}-\bm{x}^{*}$. At the same time, the learnable $\bm{h}^{*}$ will replace $\bm{h}$ and be transmitted to the decoder for high-quality reconstruction. In general, this is a special supervised learning that enables the rapid acquisition of SSI while leveraging the generalization capability of neural networks to overcome estimation errors introduced by the algorithm. Besides, the PTL-based masking strategy enjoys the advantages of negligible training overload due to the differentiable linear layer design and multi-viewpoints awareness, namely we can input multiple $\bm{h}$ into the networks for training at one time to realize a better generalization.
@@ III-C Entropy model-based transmission
For the reliable and efficient transmission, we provide a rate-adaptive JSC codec based on hyperprior-assisted non-linear transform. The reason why rate adaption is necessary lies on the masks of irregular sizes, which yield an inefficient transmission if we send images of different masked regions with a same output length. Fortunately, a promising solution is provided by [28] and [29] who uses the entropy model for image transmission with the assistance of latent hyperprior. Therefore, we provide a CNN-based autoencoder for rate-adaptive transmission by incorporating channel coding into the entropy model based framework. The architecture can be found in Fig. 7. Specifically, the masked image $\bar{\bm{x}}$ is first converted to features $\bm{v}$ via non-linear transform composed of convolutional layer and normalization. After that, spatial dependencies is captured by analysis transform on features, and we model the feature latent as scale Gaussian variables via the hyperprior $\bm{s}$, which enables a more precise computation of entropy on each component. Then, the components of features select FC layers of different output dimensions according to the respective computed entropy, and hence realizes an effective rate adaption transmission. At the receiver, the reverse operations are conducted by synthesis transform and inverse non-linear transform modules, and finally the model outputs a reconstructed masked image $\tilde{\bm{x}}$. For the training, the rate-distortion trade-off is utilized as the loss function, which is based on a variational inference, i.e.
@@
The detailed computation can be found in [29]. Note that $\bm{v}$ and $\bm{s}$ are both functions of $\bar{\bm{x}}$, and $a$ denotes the trade-off parameter between rate and distortions. In summary, the integration of adaptive mask generation with an entropy model-based codec achieves both SSI robustness and low-rate transmission. Due to above design, we find that even though RWZC encounters an unfamiliar samples during the inference process, the learning of the distribution with the homography adaption will allow the framework to effectively masking and transmitting $\bm{x}$, which will be shown in the cross-dataset testing later.